# Hybrid Beamforming Design for Near-Field ISAC with Modular XL-MIMO

### v1

@@ Abstract
A novel modular extremely large-scale multiple-input-multiple-output (XL-MIMO) integrated sensing and communication (ISAC) framework is proposed in this paper. We consider a downlink ISAC scenario and exploit the modular array architecture to enhance the communication spectral efficiency and sensing resolution while reducing the channel modeling complexity by employing the hybrid spherical and planar wavefront model. Considering the hybrid digital-analog structure inherent to modular arrays, we formulate a joint analog-digital beamforming design problem based on the communication spectral efficiency and sensing signal-to-clutter-plus-noise ratio (SCNR). By exploring the structural similarity of the communication and sensing channels, it is proved that the optimal transmit covariance matrix lies in the subspace spanned by the subarray response vectors, yielding a closed-form solution for the optimal analog beamformer. Consequently, the joint design problem is transformed into a low-dimensional rank-constrained digital beamformer optimization. We first propose a manifold optimization method that directly optimizes the digital beamformer on the rank-constrained Stiefel manifold. Additionally, we develop an semidefinite relaxation (SDR)-based approach that relaxes the rank constraint and employ the randomization technique to obtain a near-optimal solution. Simulation results demonstrate the effectiveness of the proposed modular XL-MIMO ISAC framework and algorithms, also providing valuable insights: i) The widely-spaced modular array achieves higher communication spectral efficiency than the collocated array under the same SCNR threshold; ii) The SDR-based algorithm achieves better performance, while the manifold optimization method offers lower computational complexity; iii) Key system parameters, such as the number of RF chains, subarray scale, non-line-of-sigh paths, and communication-sensing channel correlation, significantly impact the ISAC performance.
@@ I Introduction
The sixth generation wireless systems (6G) have been envisioned as a vital enabler for numerous emerging applications, such as intelligent manufacturing and smart transportation[1, 2]. The challenging problem is to satisfy the requirements of these applications for high-capacity communications and high-resolution sensing, which motivates the development of integrated sensing and communications (ISAC) technologies. The advant of extremely large-scale multiple-input and multiple-output (XL-MIMO) and the exploration of millimeter wave (mmWave)/sub- terahertz (THz) frequency bands in 6G systems lead to a growing convergence between communication and sensing in terms of channel characteristics and signal processing techniques. This convergence makes it feasible to achieve high-precision sensing and high-speed communication simultaneously using integrated waveforms and hardware platforms, offering several advantages such as reduced hardware costs, improved spectral efficiency, and mutual benefits between sensing and communication functionalities.
@@
However, the large array aperture and high operating frequencies adopted in 6G systems significantly extend the Rayleigh distance, making it more likely for communication users and sensing targets to reside in the near-field region[3]. The near-field effect not only invalidates the traditional far-field assumption-based propagation channel models but also introduces new opportunities and challenges for ISAC techniques. The work in [4] revealed that the near-field effect can potentially enhance both communication and sensing performance. However, [4] also shows that the near-field effect introduces new coupling between the two functionalities, necessitating efficient near-field beamforming designs to strike a performance balance. In [5], the authors proposed a near-field ISAC framework that exploits the additional distance dimension for optimal waveform design, showcasing performance gains over conventional far-field ISAC systems. Building upon this, [6] proposed an efficient iterative near-field beamforming algorithm for multi-target detection, demonstrating significant improvements in localization accuracy compared to traditional far-field techniques. Furthermore, the authors in [7] conducted a comprehensive performance analysis for near-field ISAC systems in both uplink and downlink scenarios, characterizing the achievable performance regions under sensing-centric, communications-centric, and Pareto optimal designs. Despite these advancements, the high model and computational complexity, along with the substantial hardware deployment costs associated with near-field XL-MIMO ISAC systems, pose significant challenges for practical implementation in realistic scenarios[4].
@@
To address these challenges, a novel modular XL-MIMO architecture, also known as widely-spaced multi-subarray (WSMS), has recently emerged as a promising solution[8, 9, 10, 11, 12]. This architecture consists of multiple modular subarrays, each composed of a flexible number of array elements with typical half-wavelength spacing, while the subarrays themselves are separated by relatively large distances[9]. The modular XL-MIMO architecture offers several advantages, such as reduced hardware costs, lower power consumption, and ease of flexible deployment in practical scenarios[12]. Moreover, the enlarged array aperture resulting from the wide spacing between subarrays exhibits a larger near-field region, potentially enhancing the performance of near-field ISAC systems. The studies have demonstrated that modular XL-MIMO, compared to collocated arrays, exhibits a more pronounced near-field effect[9], provides both inter-path and intra-path multiplexing gains to improve spectrum efficiency[8], and better adapts to the spatial non-stationarity of the channel[13]. Furthermore, the pronounced near-field effect of modular XL-MIMO also offers significant advantages for sensing. In [12], the authors investigated the potential of near-field localization with modular XL-MIMO by analyzing the Cramér-Rao Bound (CRB) for angle and range estimation, revealing that the increased array aperture and angular span of the modular array significantly enhance the near-field localization performance compared to traditional collocated arrays.
Despite the benefits of modular XL-MIMO for both near-field communication and sensing, the design and optimization of modular XL-MIMO ISAC systems remain unexplored. One of the major challenges lies in the beamforming design for downlink ISAC scenarios, where the partially-connected hybrid digital and analog architecture imposes an additional block-diagonal constraint on the analog beamformer, leading to an intractable non-convex optimization problem. The authors in [14] proposed a two-stage hybrid beamforming (HBF) method, where the analog stage computes the contribution of each submatrix on the diagonal elements of the analog beamformer to the communication spectral efficiency, and an alternating optimization algorithm is employed to solve the problem. However, this approach is not applicable to ISAC systems, as the coupled performance of communication and sensing makes it complex and impractical to separately quantify the contribution of each diagonal submatrix. For HBF design in traditional far-field MIMO-ISAC systems, [15] proposed a manifold optimization-based alternating algorithm to directly optimize the analog and digital beamformers in an iterative manner. However, it becomes inefficient and unsuitable for modular XL-MIMO ISAC systems due to the invalidity of the planar-wave assumption and the high computational complexity caused by the massive number of antennas. Moreover, the diverse designs for ISAC beamforming, such as communication-centric, sensing-centric, and Pareto-optimal approaches, often require different analog beamforming algorithms, further increasing the complexity of the problem.
@@
To fill up the research gap and address the aforementioned challenges, we propose a low-complexity hybrid beamforming design tailored for modular XL-MIMO ISAC systems in this paper. The main contributions of the paper are summarized as follows:
@@
  • We propose a novel modular XL-MIMO ISAC framework considering a downlink sensing scenario, where a base station (BS) equipped with modular XL-arrays serves a multi-antenna communication user while sensing a target in the presence of multiple interferences. Considering the relatively small size of the subarrays compared to the entire array, we employ the hybrid spherical and planar wavefront model (HSPM) to characterize the communication and sensing channels, as the user and target are more likely located in the far-field of each subarray but the near-field of the entire array. The HSPM greatly simplifies channel modeling, striking a balance between model complexity and accuracy.
@@
  • We formulate a joint analog-digital beamforming optimization problem based on communication spectral efficiency and sensing signal-to-clutter-plus-noise ratio (SCNR). By exploiting the structural similarity between the communication and sensing HSPM channels, we represent both channels using subarray response vectors. We then prove that the optimal transmit convariance matrix lies in the subspace spanned by these subarray response vectors, leading to a closed-form solution for the optimal analog beamforming matrix. Consequently, the original joint-design problem is transformed into a lower-dimensional digital beamforming optimization problem, significantly reducing the complexity.
@@
  • As the number of transmit data streams is limited by the user antennas, the digital beamforming optimization typically becomes rank-constrained and non-convex. We first propose a manifold-based algorithm that directly optimizes the digital beamformer on the rank-constrained space. Specifically, we derive the semi-closed form of the optimal digital beamformer, which forms a complex Stiefel mainifold. Then, we employ a Riemannian gradient descent approach with logarithmic barrier functions to obtain the local optimum. Despite the efficiency of the manifold optimization algorithm, its performance may be sensitive to initialization. To mitigate this issue, we develop a semidefinite relaxation (SDR)-based approach that relaxes the non-convex rank constraint and transforms the problem into a semidefinite programming (SDP) one, which can be solved to obtain a near-optimal solution through the randomization technique.
  • We conduct extensive simulations to validate the effectiveness of the proposed modular XL-MIMO ISAC framework and algorithms. The results reveal several key findings: i) The widely-spaced modular array achieves higher communication spectral efficiency compared to the traditional collocated array under the same array aperture and sensing SCNR threshold. ii) The SDR-based algorithm demonstrates superior performance in terms of solution quality, while the manifold-based algorithm offers faster convergence and lower complexity. iii) As the number of non-line-of-sight (NLoS) paths increases, both algorithms achieve higher communication spectral efficiencys under the same SCNR threshold. iv) Other system parameters, such as the number of RF chains, subarray scale, user distance, and the correlation between communication and sensing channels, significantly impact the ISAC performance. These insights provide valuable design guidelines for modular XL-MIMO ISAC systems.
@@
The rest of this paper is organized as follows. Section II introduces the system model, HSPM, and transmmitted signal model. Section III introduces the channel models for communication and sensing, along with their respective performance metrics, and presents the problem formulation. Section IV proposes two low-complexity algorithms, namely the SDR-based and manifold-based approaches, to solve the optimization problem. Simulation results are presented in Section V to validate the effectiveness of the proposed framework and algorithms, followed by concluding remarks in Section VI.
@@
Notations: We use boldface lower-case and upper-caseletters to denote column vectors and matrices, respectively. $\mathbb{C}$ denotes the set of complex number and $(\cdot)^{T}$, $(\cdot)^{H}$, $(\cdot)^{*}$ , $(\cdot)^{-1}$, $(\cdot)^{\dagger}$ denote the transposition, conjugate transposition, conjugate, inverse and pseudo-inverse, respectively. $\mathbb{E}(\cdot)$ denotes expectation. $\otimes$ denotes the Kronecker product. $\mathbf{A}\succeq\mathbf{0}$ indicates that the matrix $\mathbf{A}$ is positive semi-definite. $\mathbf{I}_{M}$ indicates an $M\times M$ identity matrix. $\det(\cdot)$ and $\text{tr}(\cdot)$ denote the determinant and trace of a matrix, respectively.
@@ II System Model
As illustrated in Fig. 1, we consider a downlink ISAC system, where the ISAC BS equipped with modular XL-arrays communicates with a multi-antenna user while simultaneously sensing a radar target. To suppress leakage signals from transmitter and receive the clear sensing echoes, the BS is assumed to be equipped with two spatially widely separated XL-arrays. The total number of transmit and receive antennas is $N=KM$, with $K$ being the number of subarrays and $M$ being the number of antenna elements within each subarray. The communication user has $N_{c}$ antennas, with $N_{c}\ll N$.
The antenna spacing within each subarray is $d=\frac{\lambda}{2}$, where $\lambda$ is the wavelength. For each subarray, we select its center element as the reference antenna. We denote the inter-subarray spacing $d_{s}=\Gamma d$ as the distance between the reference antennas of the adjacent subarrays, where $\Gamma\geq M$. There are two main reasons for considering $d_{s}$ to be much larger than $Md$. Firstly, this is necessary to accommodate practical mounting structures, such as modular XL-arrays mounted on building facades that are separated by windows. Secondly, a large inter-subarray spacing can result in an expanded array aperture, denoted by $S$, which in turn increases the near-field range of the overall antenna array, i.e., $\frac{2S^{2}}{\lambda}$. This expansion enables user who was initially located in the far field of a collocated array to now be within the near-field range of the modular array, thereby allowing them to benefit from various advantages.
The transmit and receive XL-arrays of the BS are placed along the $x$-axis, symmetrically centered around the origin. Due to the wide separation of the receive and transmit arrays, we denote the distance from the reference antenna of the first subarray to the origin as $D_{0}$. Thus, the position of $m$-th array element of $k$-th subarray at Tx, where $m\in\mathcal{M}\triangleq\{1,2,\ldots,M\}$ and $k\in\mathcal{K}\triangleq\{1,2,\ldots,K\}$, can be represented as $\mathbf{l}_{k,m}^{t}=({x}_{k,m}^{t},0)$, with ${x}_{k,m}^{t}=D_{0}+(k-1)d_{s}+(m-1)d$. Similarly, the position of $m$-th array element of $k$-th subarray at Rx is $\mathbf{l}_{k,m}^{r}=({x}_{k,m}^{r},0)$, where ${x}_{k,m}^{r}=-D_{0}-(k-1)d_{s}-(m-1)d$. Suppose that a user, target, scatterer or interference is located at $\mathbf{l}_{q}=(r\sin\theta,r\cos\theta)$, where $r$ is its distance from the origin, and $\theta\in\left[-\frac{\pi}{2},\frac{\pi}{2}\right]$ is its angle with respect to the positive $y$-axis.
@@
II-A Near Field HPSW Model for Modular XL-array
Based on the relationship between $r$ and the Rayleigh distance $\frac{2S^{2}}{\lambda}$, the space can be roughly partitioned into three parts, each corresponding to a specific model, namely, the planar-wave model (PWM), spherical-wave model (SWM), and HSPM, as elaborated below.
II-A1 Conventional PWM
When $r\geq\frac{2S^{2}}{\lambda}$, the user, target, scatterer or interference is located in the far-field region of the entire array at the Tx/Rx of the BS. In this case, the PWM is suitable for far-field propagation region, where the transmit and receive array response vectors of the BS can be respectively denoted as:
$\mathbf{a}_{t}^{\text{PWM}}\left(\theta\right)={{\left[{{e}^{-j\frac{2\pi}{\lambda}\varphi_{k,m}^{t}}}\right]}_{\forall k\in\mathcal{K},\forall m\in\mathcal{M}}^{T}},$ (1)
$\mathbf{a}_{r}^{\text{PWM}}\left(\theta\right)={{\left[{{e}^{-j\frac{2\pi}{\lambda}\varphi_{k,m}^{r}}}\right]}_{\forall k\in\mathcal{K},\forall m\in\mathcal{M}}^{T}},$
where $\varphi_{k,m}^{t}=x_{k,m}^{t}sin\theta$ and $\varphi_{k,m}^{r}=x_{k,m}^{r}sin\theta$ denote the virtual angles.
II-A2 Accurate SWM
When $r\leq\frac{2(M-1)^{2}d^{2}}{\lambda}$, the user/target/scatter is not only situated in the near-field region of the entire antenna array, but also within the near-field region of each sub-array. Consequently, the accurate SWM is applicable, leveraging the exact distances between $\mathbf{l}_{q}$ and each array element at Tx/Rx. The distances between $\mathbf{l}_{q}$ and the $m$-th array element in $k$-th subarray at Tx and Rx are respectively expressed as
$r_{k,m}^{t}=\left\|\mathbf{l}_{q}-\mathbf{l}_{k,m}^{t}\right\|=\sqrt{r^{2}-2r({x}_{k,m}^{t})\sin\theta+({x}_{k,m}^{t})^{2}},$ (2)
$r_{k,m}^{r}=\left\|\mathbf{l}_{q}-\mathbf{l}_{k,m}^{r}\right\|=\sqrt{r^{2}-2r({x}_{k,m}^{r})\sin\theta+({x}_{k,m}^{r})^{2}}.$
Therefore, the array response vectors at Tx and Rx for $\mathbf{l}_{q}$ are respectively given by[11, 10]
$\mathbf{a}_{t}^{\text{SWM}}\left(\theta\right)={{\left[{{e}^{-j\frac{2\pi}{\lambda}r_{k,m}^{t}}}\right]}_{\forall k\in\mathcal{K},\forall m\in\mathcal{M}}^{T}},$ (3)
$\mathbf{a}_{r}^{\text{SWM}}\left(\theta\right)={{\left[{{e}^{-j\frac{2\pi}{\lambda}r_{k,m}^{r}}}\right]}_{\forall k\in\mathcal{K},\forall m\in\mathcal{M}}^{T}}.$
II-A3 Harmonic HSPM
When $\frac{2(M-1)^{2}d^{2}}{\lambda}<r<\frac{2S^{2}}{\lambda}$, on one hand, SWM needs to be considered among subarrays due to $r<\frac{2S^{2}}{\lambda}$; on the other hand, PWM needs to be considered within a subarray due to $r>\frac{2(M-1)^{2}d^{2}}{\lambda}$. As such, the HSPM, combining spherical wave propagation between subarrays and plane wave propagation within subarrays is applicable. In contrast to existing research that employs uniform HSPM for common angle of arrival (AoA)s/ angle of departure (AoD)s[10], we consider the spherical wavefronts with phase variation across subarrays, and utilize HSPM for distinct AoAs/AoDs. Specifically, by representing the position of the $k$-th subarray using the location of the reference antenna in the $k$-th subarray, i.e., $\mathbf{l}_{k}^{t}=\mathbf{l}_{k,1}^{t}$ and $\mathbf{l}_{k}^{r}=\mathbf{l}_{k,1}^{r}$, the Tx and Rx array response vectors can be obtained as
$\mathbf{a}_{t}\!^{\text{HSPM}}\!\!\left(r,\theta\right)={{\left[{{e}\!^{-j\frac{2\pi}{\lambda}\left\|{{\mathbf{l}}_{q}}-\mathbf{l}_{k}^{t}\right\|}}{{e}^{-j\frac{2\pi}{\lambda}(m-1)d\sin\theta_{k}^{t}}}\right]}_{\forall k\in\mathcal{K},\forall m\in\mathcal{M}}^{T}},$ (4)
$\mathbf{a}_{r}\!^{\text{HSPM}}\!\!\left(r,\theta\right)={{\left[{{{e}\!^{-j\frac{2\pi}{\lambda}\left\|{{\mathbf{l}}_{q}}-\mathbf{l}_{k}^{r}\right\|}}}{{e}^{-j\frac{2\pi}{\lambda}(m-1)d\sin\theta_{k}^{r}}}\right]}_{\forall k\in\mathcal{K},\forall m\in\mathcal{M}}^{T}},$
where ${\theta}_{k}^{t}$ and ${\theta}_{k}^{r}$ are the angles of $\mathbf{l}_{q}$ as observed from the $k$-th subarray at Tx and Rx. Following Fig. 1, we have
$\sin\theta_{k}^{t}=\frac{r\sin\theta-{{D}_{0}}-\left(k-1\right){{d}_{s}}}{\left\|{{\mathbf{l}}_{q}}-\mathbf{l}_{k}^{t}\right\|},$ (5)
$\sin\theta_{k}^{r}=\frac{r\sin\theta+{{D}_{0}}+\left(k-1\right){{d}_{s}}}{\left\|{{\mathbf{l}}_{q}}-\mathbf{l}_{k}^{r}\right\|}.$
In conclusion, SWM accurately captures the signal amplitude and phase variations of all array elements, making it the most accurate and generic model. However, it is associated with high complexity in modeling and signal processing. The PWM serves as an approximation of the SWM when the array aperture $S$ is significantly smaller than the distance $r$. However, the approximation error of the PWM grows as the array aperture increases[16]. Therefore, in the considered modular XL-MIMO architecture, PWM becomes inaccurate. By comparison, HSPM outperforms PWM in terms of accuracy and exhibits lower complexity compared to SWM, making it a more tractable and accurate near-field array model[17, 10].
II-B Transmit ISAC Signal
Due to the extremely large number of antennas in XL-MIMO systems, the implementation cost of a fully digital beamforming architecture is unaffordable. Therefore, we consider a modular XL-MIMO hybrid digital and analog architecture, where $M_{{\text{RF}}}$ RF chains are assigned to control one subarray. The total number of RF chains at Tx is $N_{\text{RF}}=KM_{\text{RF}}$. Additionally, due to the multiple antennas deployed at the communication user, the transmission of multiple data streams $N_{s}$ for a single user can be realized, where $N_{s}\leq N_{\text{RF}}$.

### v2

# Near-Field Hybrid Beamforming Design for Modular XL-MIMO ISAC Systems

@@ Abstract
A novel modular extremely large-scale multiple-input-multiple-output (XL-MIMO) integrated sensing and communication (ISAC) system is investigated in this paper. A downlink sensing scenario is considered, where the modular XL-MIMO BS, consisting of widely-spaced identical subarrays, simultaneously serves a multi-antenna user and performs target sensing in the presence of multiple interferences. Due to the small subarray aperture and large array spacing, a piecewise-far-field channel model is employed to characterize both communication and sensing channels, where far-field propagation exists within each subarray while near-field effects dominate among subarrays. Based on the group-connected hybrid architecture, a joint analog-digital beamforming problem is formulated to optimize communication spectral efficiency while satisfying the sensing signal-to-clutter-plus-noise ratio (SCNR) requirement. Then, the optimal transmit covariance matrix is proved to lie in the subspace spanned by subarray response vectors. Based on this, a closed-form solution for the optimal analog beamformer is derived, and the original joint analog-digital optimization problem is transformed into a low-dimensional rank-constrained digital beamforming problem. Firstly, the semi-closed form of the optimal digital beamformer is derived and shown to form a complex Stiefel manifold. Based on this structure, a joint Riemannian-Euclidean gradient descent algorithm is developed for iterative optimization. Then, an semidefinite relaxation-based approach is proposed, where the near-optimal solution is obtained through rank constraint relaxation and randomization. The superiority of the proposed algorithms is extensively validated through, revealing that the optimal subarray scale balances spatial multiplexing and beamforming gains based on user distance, while increasing subarray numbers significantly enhances range resolution due to more pronounced spherical wavefronts.
@@ I Introduction
The sixth generation wireless systems (6G) have been envisioned as a vital enabler for numerous emerging applications, such as intelligent manufacturing and smart transportation[1]. The challenging problem is to satisfy the requirements of these applications for high-capacity communications and high-resolution sensing, which motivates the development of integrated sensing and communications (ISAC) technologies[2]. The advant of extremely large-scale multiple-input and multiple-output (XL-MIMO) and the exploration of millimeter wave (mmWave)/sub- terahertz (THz) frequency bands in 6G systems lead to a growing convergence between communication and sensing in terms of channel characteristics and signal processing techniques. This convergence makes it feasible to achieve high-precision sensing and high-speed communication simultaneously using integrated waveforms and hardware platforms, offering several advantages such as reduced hardware costs, improved spectral efficiency, and mutual benefits between sensing and communication functionalities[3, 4].
@@
However, the large array aperture and high operating frequencies adopted in 6G systems significantly extend the Rayleigh distance, making the near-field effect more prominent for communication users and sensing targets[5]. Recent studies have shown that near-field propagation not only invalidates traditional far-field channel models but also introduces new opportunities and challenges for ISAC techniques[6, 3, 7, 8, 9]. The work in [7] revealed that the near-field effect can potentially enhance both communication and sensing performance. The authors in [8] proposed a novel near-field ISAC framework that leverages the additional distance dimension for optimal waveform design, demonstrating superior performance compared to traditional far-field ISAC systems. Subsequently, [9] proposed an efficient iterative near-field beamforming algorithm for multi-target detection, achieving significant enhancements in localization accuracy over conventional far-field techniques. Despite these advancements, the high computational complexity and substantial hardware deployment costs associated with near-field XL-MIMO ISAC systems pose significant challenges for practical implementation in realistic scenarios [7].
@@
To address these challenges, a novel modular XL-MIMO architecture, also known as widely-spaced multi-subarray (WSMS), has recently emerged as a promising solution[10, 11, 12, 13, 14]. This architecture consists of multiple modular subarrays, each composed of a flexible number of array elements with typical half-wavelength spacing, while the subarrays are separated by relatively large distances[11]. The modular XL-MIMO architecture offers advantages such as reduced hardware costs, lower power consumption, and ease of flexible deployment in practical scenarios [14]. Moreover, the enlarged array aperture resulting from the wide spacing between subarrays exhibits a larger near-field region, potentially enhancing the performance of near-field ISAC systems. Studies have shown that modular XL-MIMO, compared to collocated arrays, exhibits a more pronounced near-field effect [11], provides both inter-path and intra-path multiplexing gains to improve spectral efficiency[10], and better adapts to the spatial non-stationarity of the channel[15]. In [14], the authors investigated the potential of near-field localization with modular XL-MIMO by analyzing the Cramér-Rao Bound (CRB) for angle and range estimation, revealing that the increased array aperture and angular span of the modular array significantly enhance the near-field localization performance compared to traditional collocated arrays.
Despite the benefits of modular XL-MIMO for both near-field communication and sensing, the design and optimization of modular XL-MIMO ISAC systems remain unexplored. The primary challenge lies in the hybrid beamforming design, where the group-connected hybrid architecture imposes a block-diagonal constraint on the analog beamformer, resulting in an intractable non-convex optimization problem[16]. Existing hybrid beamforming (HBF) methods [17, 18] mainly focus on optimizing for a single communication functionality and cannot effectively handle the block-diagonal constraint in modular XL-MIMO ISAC systems, where the coupled communication and sensing performance lead to potential conflicts and make it impractical to separately optimize each diagonal submatrix or directly apply iterative optimization methods. Moreover, the extremely large antenna arrays introduce high computational complexity, making conventional optimization algorithms impractical for real-world implementation. Additionally, the modular structure exhibits unique channel propagation characteristics, with the entire large-scale array operating in the near-field while the small-scale subarrays are in the far-field[19]. This hybrid propagation characteristic presents an opportunity for complexity reduction, as it naturally decomposes into far-field propagation within each subarray and near-field effects between subarrays, mitigating the nonlinear phase variation of received signals[20]. Therefore, it is necessary to develop an efficient optimization algorithm that exploits the hybrid propagation channel structure while addressing both the structural constraints and performance requirements.
@@
In this paper, we propose a low-complexity hybrid beamforming design tailored for modular XL-MIMO ISAC systems. The main contributions of the paper are summarized as follows:
@@
  • We consider a downlink ISAC scenario, where a base station (BS), equipped with modular extremely large-scale (XL)-arrays and the hybrid digital and analog (HDA) architecture, serves a multi-antenna communication user while sensing a target in the presence of multiple interferences. The number of active radio frequency (RF) chains is adapted to the communication and sensing channel conditions to provide sufficient spatial degrees of freedom (DoF)s. Considering the relatively small size of the subarrays compared to the entire array, we employ the piecewise-far-field channel model to characterize the communication and sensing channels, as the user and target are typically located in the far-field of each subarray but the near-field of the entire array.
@@
  • We formulate a joint analog-digital beamforming optimization problem based on communication spectral efficiency and sensing signal-to-clutter-plus-noise ratio (SCNR). Then, we derive a closed-form solution for the optimal analog beamformer by proving that the optimal transmit covariance matrix lies in the subspace spanned by the communication and sensing subarray response vectors. This theoretical insight transforms the original joint analog-digital optimization problem into a low-dimensional digital beamforming problem, substantially reducing the computational complexity.
@@
  • We develop two efficient algorithms to solve the rank-constrained digital beamforming optimization. First, we derive the semi-closed form of the optimal digital beamformer, which forms a complex Stiefel mainifold. Based on this, we propose a Riemannian-Euclidean joint gradient descent algorithm to iteratively obtain the local optimal solution. Additionally, we develop a semidefinite relaxation (SDR)-based approach that relaxes the non-convex rank constraint and transforms the problem into a semidefinite programming (SDP) one, which can be solved to obtain a near-optimal solution through the randomization technique.
  • We conduct extensive simulations to validate the effectiveness of the proposed modular XL-MIMO ISAC algorithms. The results reveal several key findings: i) The proposed SDR-based algorithm achieves superior performance over conventional HBF algorithms. ii) The optimal subarray scale varies with user distance, representing a tradeoff between spatial multiplexing and beamforming gains. iii) Increasing the number of subarrays significantly enhances the range resolution due to more pronounced spherical wavefronts across subarrays. These insights provide valuable design guidelines for modular XL-MIMO ISAC systems.
@@
The rest of this paper is organized as follows. Section II introduces the system settings and the transmmitted signal model. Section III introduces communication and sensing channel models, along with their respective performance metrics, and presents the problem formulation. Section IV proposes two low-complexity HBF algorithms. Simulation results are presented in Section V, followed by concluding remarks in Section VI.
@@
Notations: We use boldface lower-case and upper-case letters to denote column vectors and matrices, respectively. $\mathbb{C}$ denotes the set of complex number and $(\cdot)^{T}$, $(\cdot)^{H}$, $(\cdot)^{*}$ , $(\cdot)^{-1}$, $(\cdot)^{\dagger}$ denote the transposition, conjugate transposition, conjugate, inverse and pseudo-inverse, respectively. $\mathbb{E}(\cdot)$ denotes expectation. $\otimes$ denotes the Kronecker product. $\mathbf{A}\succeq\mathbf{0}$ indicates that the matrix $\mathbf{A}$ is positive semi-definite. $\mathbf{I}_{M}$ indicates an $M\times M$ identity matrix. The $\mathrm{diag}(\cdot)$ operator creates a diagonal matrix with the input elements placed along the main diagonal, while all other entries are set to zero. $\det(\cdot)$ and $\text{tr}(\cdot)$ denote the determinant and trace of a matrix, respectively.
@@ II System Model
As illustrated in Fig. 1, we consider a downlink ISAC system, where the ISAC BS equipped with transmit and receive modular XL-arrays communicates with a multi-antenna user while simultaneously sensing a radar target. Both transmit and receive arrays are deployed along the $x$-axis, symmetrically centered around the origin, and widely separated in space to suppress signal leakage from transmitter and receive clear sensing echoes[21]. The total number of transmit and receive antennas is $N=KM$, with $K$ subarrays and $M$ antenna elements within each subarray. The communication user has $N_{c}$ antennas, where $N_{c}\ll N$.
The antenna spacing within each subarray is $d=\frac{\lambda}{2}$, where $\lambda$ is the wavelength. For each subarray, we select its first element as the reference antenna. We denote the inter-subarray spacing $d_{s}=\Gamma d$ as the distance between the reference antennas of the adjacent subarrays, where $\Gamma\geq M$. There are two main reasons for considering $d_{s}$ to be much larger than $Md$. Firstly, it accommodates practical mounting structures, such as modular XL-arrays mounted on building facades that are separated by windows[12]. Secondly, a large inter-subarray spacing expands the array aperture, $S$, which increases the near-field range of the overall antenna array to $\frac{2S^{2}}{\lambda}$. This expansion allows users initially located in the far-field of a collocated array to be within the near-field range of the modular array, enabling them to benefit from various advantages.
Assuming the distance between the transmit and receive arrays is $2D_{0}$, the position of the $m$-th element in the $k$-th subarray at transmitter (Tx) and receiver (Rx) is given by $\mathbf{l}_{k,m}^{t}=({x}_{k,m}^{t},0)$ and $\mathbf{l}_{k,m}^{r}=({x}_{k,m}^{r},0)$, respectively, where ${x}_{k,m}^{t}={D_{0}}+(k-1)d_{s}+(m-1)d$ and ${x}_{k,m}^{r}=-{D_{0}}-(k-1)d_{s}-(m-1)d$, with $m\in\mathcal{M}\triangleq\{1,2,\ldots,M\}$ and $k\in\mathcal{K}\triangleq\{1,2,\ldots,K\}$. Suppose that a user, target, scatterer or interference located at $\mathbf{l}_{q}=(r\sin\theta,r\cos\theta)$is characterized by its distance $r$ from the origin and angle $\theta\in\left[-\frac{\pi}{2},\frac{\pi}{2}\right]$ with respect to the positive $y$-axis.
@@
II-A Transmit ISAC Signal
As illustrated in Fig. 2, we consider a group-connected HDA architecture, where each subarray is equipped with dedicated RF chains that are connected to all antennas within that subarray but cannot be shared across subarrays due to the the hardware limitations[10, 16]. Specifically, each subarray is equipped with $\tilde{M}_{\text{RF}}$ RF chains, where each chain is connected to all antennas within that subarray, resulting in a total of $\tilde{N}_{\text{RF}}=K\tilde{M}_{\text{RF}}$ available RF chains at the Tx. For energy efficiency, digital switches are employed after the digital precoder to adaptively control the number of active RF chains according to the practical communication and sensing channel conditions [22]. We denote the number of active RF chains as $M_{\text{RF}}$ per subarray, resulting in a total of $N_{\text{RF}}=KM_{\text{RF}}$ active RF chains. For data transmission, the multi-antenna user can can support the transmission of $N_{s}$ data streams, where $N_{s}\leq N_{\text{RF}}$.